OpenStudy (anonymous):

Challenge question: -7=1 x - 5

OpenStudy (texaschic101):

is the problem -7 = 1/5x ??

OpenStudy (anonymous):

x=-35 if the equation is -7=1 (x/5)

OpenStudy (texaschic101):

that would be the same answer as if the equation was -7 = 1/5x. Because you would multiply the entire equation by 5, then solve for x

OpenStudy (anonymous):

(texaschic) yes :]

OpenStudy (texaschic101):

-7 = 1/5x (multiply the equation by 5) -35 = x

OpenStudy (texaschic101):

or you could divide both sides by 1/5 -7 = 1/5x (divide by 1/5) -7/(1/5) = x -7 * 5/1 = x -35/1 or just -35 = x

OpenStudy (anonymous):

Thanks:)

OpenStudy (texaschic101):

no problem :)